OpenStudy (anonymous):

Find the slope of the line that goes through the points (12,3) and (13,4).Enter your answer as an integer or a reduced fraction in the form A/B

OpenStudy (anonymous):

since slope of a line is =(y2-y1)/(x2-x1) point (12,3)where x1=12 and y1=3and the other point (13,4) where x2=13 and y2=4 substitutingthe values into the formular we have slope=(4-3)/(13-12) slope=1/1 since we were told to reduce it in the form A/B THEREFORE your answer is slope=1/1